OpenStudy (anonymous):

When multiplying and simplifying (a-b) (a+b) (a+b)^2 it equals 2a^2+2B^2 right?

OpenStudy (anonymous):

(a-b)(a+b) = a^2 - b^2 (a^2-b^2)(a^2+2ab+b^2) = a^4+2a^3b - 2ab^3-b^4